# V. Distance Estimation with Neutrinos Alone

### v1

@@ Abstract
Neutrinos are pivotal signals in multi-messenger observations of supernovae. Recent advancements in the analysis method of supernova neutrinos, especially in quantitative analysis, have significantly broadened scientific possibilities. This study demonstrates the feasibility of estimating distances to supernovae using quantitative analysis techniques for supernova neutrinos. This estimation utilizes the direct relationship between the radius of a neutron star and the distance to the supernova. The radius of a neutron star is determined with an approximate uncertainty of 10% through observations such as X-rays and gravitational waves. By integrating this information, the distance to the supernova can be estimated with an uncertainty of within 15% at a 95% confidence level. It has been established that neutrinos can pinpoint the direction of supernovae, and when combined with distance estimates, three-dimensional localization becomes achievable. This capability is vital for follow-up observations using multi-messenger approaches. Moreover, more precise distance determinations to supernovae through follow-up observations, such as optical observations, allow for accurate measurements of neutron star radii. This data, via the neutron star mass-radius relationship, could provide various insights into nuclear physics.
Supernova neutrinos (1666); Neutrino astronomy (1100); Neutrino telescopes (1100); Core-collapse supernovae (304); Neutron stars (1108);
@@ 1 Introduction
A supernova (SN) produces a large number ($\sim 10^{58}$) of neutrinos. These neutrinos, originating from thermal processes, are emitted almost isotropically. As a result, neutrinos generated in nearby supernovae (SNe) ensure their detectability (see Janka, 2017; Horiuchi & Kneller, 2018; Müller, 2019; Vartanyan & Burrows, 2023; Yamada et al., 2024, and references therein). It is noteworthy that neutrinos are emitted before electromagnetic radiation, as their production occurs while the shock wave is still confined within the star (Kistler et al., 2013; Abe et al., 2016). Therefore, observing SN neutrinos is an initial step for further multiband observational studies that utilize the broad electromagnetic wavelength, extending from radio to gamma rays. Within the realm of multi-messenger studies in SN research, neutrinos are deemed paramount from diverse scientific perspectives.
@@
In this paper, we investigate the potential of determining the distance to an SN through quantitative analysis of SN neutrinos. A pivotal element of this study is the direct relationship between the distance of the SN and the radius of the PNS. Essentially, determining the radius of the PNS facilitates the estimation of the distance to the SN, and vice versa. This study evaluates the precision with which the distance to an SN can be measured using NS radii derived from alternative observations such as X-rays.

### v2

@@ Abstract
Neutrinos are pivotal signals in multi-messenger observations of supernovae (SNe). Recent advancements in the analysis method of supernova (SN) neutrinos, especially in quantitative analysis, have significantly broadened scientific possibilities. This study demonstrates the feasibility of estimating distances to SNe using neutrinos. This estimation utilizes the direct relationship between the radius of a neutron star (NS) and the distance to the SN, which is analogous to main-sequence fitting. The radius of an NS is determined with an approximate uncertainty of 10% through observations such as X-rays and gravitational waves. By integrating this information, the distance to the SN can be estimated with an uncertainty of within 15% at a 95% confidence level. It has been established that neutrinos can pinpoint the direction of SNe, and when combined with distance estimates, three-dimensional localization becomes achievable. This capability is vital for follow-up observations using multi-messenger approaches. Moreover, more precise distance determinations to SNe through follow-up observations, such as optical observations, allow for accurate measurements of NS radii. This data, via the NS mass-radius relationship, could provide various insights into nuclear physics.
Core-collapse supernovae (304); Neutrino astronomy (1100); High energy astrophysics (739); Neutron stars (1108)
@@ 1 Introduction
A supernova (SN) produces a large number ($\sim 10^{58}$) of neutrinos. These neutrinos, originating from thermal processes, are emitted almost isotropically. As a result, neutrinos generated in nearby supernovae (SNe) ensure their detectability (see Janka, 2017; Horiuchi & Kneller, 2018; Müller, 2019; Vartanyan & Burrows, 2023; Yamada et al., 2024, and references therein). It is noteworthy that neutrinos are emitted before electromagnetic radiation, as their production occurs while the shock wave is still confined within the star (Kistler et al., 2013; Abe et al., 2016). Therefore, observing SN neutrinos is an initial step for further multiband observational studies that utilize the broad electromagnetic wavelength, extending from radio to gamma rays. In multi-messenger studies of SN research, neutrinos are regarded as extremely important from various scientific perspectives.
@@
In this paper, we investigate the potential of determining the distance to an SN through a quantitative analysis of SN neutrinos, emphasizing the straightforwardly derived relationship between the SN distance and the radius of the PNS. By determining the PNS radius, we can estimate the SN distance, and vice versa. Based on this relationship, we introduce a neutrino-based distance estimation method analogous to traditional astronomical techniques such as main-sequence fitting or spectroscopic parallax. In this method, the distance of a star is determined by comparing its absolute magnitude, derived from its position on the Hertzsprung-Russell diagram, with its observed magnitude. Similarly, our approach utilizes the thermal distribution of neutrinos, applying the Stefan-Boltzmann law and the radius of the neutron star–obtained from other observations like X-rays–to estimate distances. Thus, the simplicity of Eqs. (5) and (11) provides a straightforward quantitative framework for neutrino-based SN distance determination, analogous to how the main-sequence fitting relies on well-characterized stellar properties.
@@ 5 Summary and discussion
We simplify the analysis by focusing on late-phase neutrino emission. Specifically, neutrinos are emitted through a quasi-static diffusion process within a time-stationary neutron star density distribution. As demonstrated by Suwa et al. (2021), neutrino emission can be expressed analytically, and this study is based on that analytic formula. However, additional physical factors that may influence neutrino emission should be accounted for as systematic errors. For instance, the effect of neutrino oscillations causes flavor transitions, while the neutrino spectrum becomes nearly flavor-independent at late times. These factors include uncertainties in numerical simulations of supernovae and detector response, which are essential for a more accurate evaluation. While these factors fall beyond the primary scope of this study, their inclusion in uncertainty evaluations will be essential for a more comprehensive analysis. Consequently, the uncertainties reported here reflect only statistical errors. We are working on these aspects and systematic error evaluations, which will be reported in forthcoming papers.